# Conceptualization, Methodology, Software, Analysis, Writing and reviewing manuscript

### v1

@@ Abstract
Modeling viscoelastic behavior is crucial in engineering and biomechanics, where materials undergo time-dependent deformations, including stress relaxation, creep buckling and biological tissue development. Traditional numerical methods, like the finite element method, often require explicit meshing, artificial perturbations or embedding customised programs to capture these phenomena, adding computational complexity. In this study, we develop an energy-based physics-informed neural network (PINN) framework using an incremental approach to model viscoelastic creep, stress relaxation, buckling, and growth-induced morphogenesis. Physics consistency is ensured by training neural networks to minimize the system’s potential energy functional, implicitly satisfying equilibrium and constitutive laws. We demonstrate that this framework can naturally capture creep buckling without pre-imposed imperfections, leveraging inherent training dynamics to trigger instabilities. Furthermore, we extend our framework to biological tissue growth and morphogenesis, predicting both uniform expansion and differential growth-induced buckling in cylindrical structures. Results show that the energy-based PINN effectively predicts viscoelastic instabilities, post-buckling evolution and tissue morphological evolution, offering a promising alternative to traditional methods. This study demonstrates that PINN can be a flexible, robust tool for modeling complex, time-dependent material behavior, opening up possible applications in structural engineering, soft materials, and tissue development.
@@ 1 Introduction
Many materials, including polymers, biological tissues and soft composites, exhibit a combination of elastic energy storage and viscous energy dissipation, i.e., the viscoelasticity (Fig.1a). Understanding and predicting time-dependent phenomena are essential for numerous engineering and biological applications, such as designing advanced materials [1], optimizing structural performance [2, 3], and studying biological tissue growth and morphogenesis [4]. The complex viscoelastic behaviors, particularly in cases involving large deformations, nonlinear responses, and evolving geometries, present persistent modeling challenges.
@@
Over the past decades, computational approaches such as the finite element method (FEM) and finite difference method have been widely used to simulate viscoelastic responses, including creep, stress relaxation, buckling instabilities, and wave propagation [5, 6, 7, 8]. While these methods have been successful in many scenarios, they often require refined meshing strategies and explicit perturbations for specific problems. For example, in instability analysis, traditional FEM always needs artificial perturbations or imperfections which can lead to different buckling modes [9]. Moreover, in highly nonlinear and evolving systems, such as biological growth and morphogenesis, these conventional approaches can struggle to handle very complex geometries or require complex implementation strategies [10, 11]. The need for a more adaptive computational framework is evident.
@@
Figure 1: Viscoelastic materials and models. (a) Viscoelasticity is very common among different materials, e.g., hydrogels, bitumen [12], and soft biological tissues. (b-c) The viscoelastic models used for mechanical modeling. (b) The general Maxwell model. (c) The standard three-parameter model.
@@
Physics-informed neural networks (PINNs) have recently emerged as a promising alternative to solve some numerical problems [13]. PINNs offer a mesh-free approach that combines data and physical laws to solve differential equations by embedding the governing physics into the loss function of neural networks [13, 14]. They have been successfully applied to forward and inverse problems in fluid dynamics [15, 16], solid mechanics [17, 18], and material science [19, 20], demonstrating their capability to solve physical problems. In viscoelastic modeling , the inverse problem has been studied using a deep neural network to obtain parametric and neural-network-based viscoelastic models from limited displacement data [21]. A proposed ViscoelasticNet using PINNs with appropriate constitutive models was applied to learn stress fields of viscoelastic fluids [22]. The deep energy method was developed to solve partial differential equations governing the deformation of hyperelastic and viscoelastic materials [23]. It leverages the variational principles of mechanics, training networks to minimize a system’s potential energy functional. These pioneering applications validate the use of PINN and related deep learning techniques for mechanical simulation and viscoelastic constitutive modeling.
@@
For instability studies, buckling analysis of Kirchhoff plates has been implemented using the deep energy method, where a loss function based on the buckling governing equations is used to predict the critical buckling load [18]. While this marks significant progress for elastic instabilities, the distinct phenomenon of creep buckling of viscoelastic structures—where instability develops over time under sustained load due to viscosity—necessitates a dedicated PINN formulation capable of handling long-term temporal integration and history-dependent material laws. Indeed, for viscoelastic structures, creep buckling is one of the important failure modes, signifying failure after a period under steady load either via sudden snapping or excessive deflection [24]. Additionally, during growth and development of soft biological tissues with viscoelastic properties, the symmetry broken and geometrical stability loss are important processes for tissue morphogenesis [25, 26, 27]. Despite the large potential of PINN, extending these approaches to capture creep buckling of viscoelastic structures, and to model the intricate interplay of mechanics and material evolution in viscoelastic growth and morphogenesis, presents difficulties and remains an open challenge.
Addressing these challenges, this study develops and applies an energy-based PINN framework for modeling viscoelastic behavior. We first demonstrate that PINNs can effectively capture viscoelastic creep and stress relaxation, simulating the time-dependent behavior of materials under sustained loading. Additionally, we highlight a key advantage of PINN for instability analysis, i.e., the intrinsic training oscillations [17]. This intrinsic capability offers a potential route to study the computational stability analysis, providing a natural pathway to discovering buckling behavior. Furthermore, we extend our framework to study viscoelastic tissue growth and morphogenesis, revealing that PINNs can capture complex morphological evolutions that are difficult to achieve with standard finite element approaches. Through comprehensive benchmarking and validation, we illustrate the strengths and limitations of PINNs in modeling viscoelastic phenomena. Our findings suggest that PINNs offer a powerful and flexible tool for capturing time-dependent material behaviors, opening new possibilities for applications in soft materials and biomechanics. The remainder of this paper is organized as follows: Section 2 introduces the mathematical formulation of viscoelasticity. Section 3 details the implementation of PINN for viscoelastic simulation, and the benchmark of creep and stress relaxation. Section 4 explores the role of training oscillations in capturing creep buckling phenomena. Section 5 extends the framework to viscoelastic growth and morphogenesis, demonstrating its potential in biological applications. Finally, Section 6 discusses the advantages, limitations, and future research directions.
@@ 2.1 Governing Equation of Nonlinear Viscoelasticity
Let $\Omega$ be the configuration of a body in the three-dimensional Euclidean space $\mathbb{E}^{3}$, and the corresponding initial configuration is $\Omega_{0}$. The deformation gradient tensor is $\mathbf{F}$, and $\mathbf{F}=\mathbf{I}+\nabla\mathbf{u}$, where $\mathbf{u}$ is the displacement vector and $\mathbf{I}$ is the identity tensor. For hyperelastic neo-Hookean materials, the strain energy density can be expressed as [28]
@@
where $\lambda$ is the Lamé’s first parameter, $G$ is the shear modulus, $J=\text{det}\left(\mathbf{F}\right)$ is the Jacobian that measures the relative volume change, $I_{1}=B_{ii}$ is the first deformation invariant. $\mathbf{B}=\mathbf{F}\cdot\mathbf{F}^{T}$ is the left Cauchy-Green deformation tensor. The Cauchy stress formula can be deduced by differentiating the strain gradient density, as
@@
Based on previous studies [29, 30], the constitutive equation of viscoelastic material under large deformation can be written as
@@
where $m(t)$ is the dimensionless relaxation function and is independent of deformation. Using the general Maxwell model (Fig. 1b), $m(t)$ can be expressed by Prony series:
@@
where $\tau_{i}=\xi_{i}/E_{i}$ are the relaxation times, $m_{\infty}+\sum_{i}m_{i}=1$. If we describe viscoelasticity by the standard solid model (Fig. 1c), it has
@@ 2.2 Stress Relaxation and Creep
For the viscoelastic stress relaxation condition, the deformation remains constant, denoted as $\mathbf{F}_{0}$. Then, the Cauchy stress evolution equation can be obtained from Eq. (6), i.e.,
@@
where $\mathbf{B}_{0}=\mathbf{F}_{0}\cdot\mathbf{F}_{0}^{T}$, $J_{0}=\det\left(\mathbf{F}_{0}\right)$. Specifically, under the small deformation and one-dimensional condition, the strain remains constant as $\varepsilon_{0}$, and the stress evolution equation reduces to $\sigma(t)=E(t)\varepsilon_{0}$ [31]. When using the standard solid model Fig. 1c, the stress is
@@
The relaxation time $\tau$ is the characteristic time it takes for stress to decay significantly in a viscoelastic material, with the initial rate being $\dot{\sigma}(0)=-(E_{1}\varepsilon_{0})/{\tau}$.
Creep is defined as the tendency of a solid material to deform permanently under the influence of constant stress. Under creep conditions, the stress remains constant, i.e., $\sigma_{0}$. The strain creep function can be written as $\mathbf{e}(t)=J(t)\sigma_{0}$, where $J(t)$ is the creep compliance function and $\mathbf{e}$ is the strain tensor. The creep compliance can be obtained by applying the Laplace transform and inverting the relaxation function. The analytical solutions may not be available under large deformation, but the creep behavior can be captured easily by numerical methods, like finite element analysis. Specifically, for the small deformed one-dimensional case, the creep process can be simulated using the standard solid model (Fig. 1c)
@@
where the retardation time $\tau_{r}$ is defined as $\tau_{r}=\tau(E_{\infty}+E_{1})/E_{\infty}$, which can characterize the creep behavior. The creep rate is
@@
If the material creeps with the initial creep rate $\dot{\varepsilon}(0)={\sigma_{0}E_{1}}/[{\tau_{r}E_{\infty}(E_{\infty}+E_{1})}]$, the duration of the creep process from $\varepsilon(0)=\sigma_{0}/(E_{1}+E_{\infty})$ to $\varepsilon(\infty)=\sigma_{0}/E_{\infty}$ is $\tau_{r}$.
@@ 2.3 Creep Buckling of Viscoelastic Structures
Creep buckling refers to a specific type of structural failure that occurs in materials subjected to a sustained compressive load over an extended period [24, 32]. As the material creeps, its effective rigidity decreases. This reduction can lower the critical buckling load, i.e., the maximum load that a structure can withstand before it buckles, leading to premature failure even when loads are lower than the instantaneous buckling thresholds.
@@
where $l$ is the length of the cantilever, $E$ is the Young’s modulus of the material and $I_{g}$ is the area moment of inertia. For the cross-section with area $A$, the critical pressure is
@@
For a linear elastic thin-walled cylindrical shell under uniform radial pressure, the critical buckling pressure can be estimated by [24]
@@
where $d$ is the cylinder diameter, and $a$ is the wall thickness. If a structure, such as a slender beam or a thin-walled cylindrical shell, exhibits viscoelastic behavior, its stiffness decreases and deformation increases over time under a constant external load. As a result, its critical buckling load becomes lower than the initial elastic buckling threshold. For creep buckling analysis, a quasi-static approach is often adopted, that is, buckling occurs at a time $t_{cr}$ when the applied pressure $p$ equals the critical elastic buckling load calculated using the current relaxed modulus $E(t_{cr})$ and Eqs. (13) or (14).
@@ 2.4 Viscoelastic Tissue Growth
Viscoelasticity is a fundamental property of biological soft tissues, characterized by their ability to exhibit both elastic and viscous behavior when subjected to deformation [33, 34]. The stress modulated growth models have been developed [35, 36, 37] since the stress-growth law proposed by Fung [38]. In this kind of growth model, the deformation gradient tensor is decomposed as (Fig. 2)
@@
where $\mathbf{F}_{a}$ is the mechanical deformation tensor and $\mathbf{F}_{g}$ is the growth tensor. The growth tensor is always written as $\mathbf{F}_{g}=g_{i}\mathbf{I}$, where $g_{i}$ is the growth ratio. The growth rate is determined by the stress state [37, 25], i.e.,
@@
where $\mathbf{b}^{g}=b^{g}_{i}\mathbf{I}$ is the biochemical force vector driving growth [39], and $k$ is a constant modulating the influence of the stress on growth. For viscoelastic tissues, the stress $\sigma$ is obtained by Eq. (6), based on the deformation gradient tensor $\mathbf{F}_{a}$, which evolves and relaxes over time. Stress relaxation slows residual stress accumulation, leading to increased growth rates [25].
@@
A critical consequence of the differential or constrained growth described by the growth tensor $\mathbf{F}_{g}$, coupled with the viscoelastic response, is the generation of internal residual stresses [40, 25]. When growth is non-uniform (e.g., faster surface growth relative to the bulk) or spatially constrained, significant compressive stresses can accumulate within the tissue over time. If these growth-induced compressive stresses reach a critical threshold, the initial configuration of the tissue may lose stability, leading to a buckling event [41, 37]. This mechanical instability provides a fundamental mechanism for morphogenesis, driving the spontaneous formation of complex patterns such as wrinkles, folds, or creases observed during development and other biological processes.

### v2

@@ Abstract
Modeling viscoelastic behavior is crucial in engineering and biomechanics, where materials undergo time-dependent deformations, including stress relaxation, creep buckling and biological tissue development. Traditional numerical methods, like the finite element method, often require explicit meshing, artificial perturbations or embedding customised programs to capture these phenomena, adding computational complexity. In this study, we develop an energy-based physics-informed neural network (PINN) framework using an incremental approach to model viscoelastic creep, stress relaxation, buckling, and growth-induced morphogenesis. Physics consistency is ensured by training neural networks to minimize the system’s potential energy functional, implicitly satisfying equilibrium and constitutive laws. We demonstrate that this framework can naturally capture creep buckling without pre-imposed imperfections, leveraging inherent training dynamics to trigger instabilities. Furthermore, we extend our framework to biological tissue growth and morphogenesis, predicting both uniform expansion and differential growth-induced buckling in cylindrical structures. Results show that the energy-based PINN effectively predicts viscoelastic instabilities, post-buckling evolution and tissue morphological evolution, offering a promising alternative to traditional methods. This study demonstrates that PINN can be a flexible robust tool for modeling complex, time-dependent material behavior, opening possible applications in structural engineering, soft materials, and tissue development.
@@ 1 Introduction
Many materials, including polymers, biological tissues and soft composites, exhibit combination of elastic energy storage and viscous energy dissipation [Review_1], i.e., the viscoelastic properties (Fig.1a). Understanding and predicting the time-dependent phenomena are essential for numerous engineering and biological applications [Review_3], such as designing advanced materials [Lin.2020], optimizing structural performance [Sun.2013, Lin.2021], and studying biological tissue growth and morphogenesis [Collinet.2021]. The complex viscoelastic behaviors, particularly in cases involving large deformations, nonlinear responses, and evolving geometries, present persistent modeling challenges.
@@
Over the past decades, computational approaches such as the finite element method (FEM) and finite difference method have been widely used to simulate viscoelastic responses, including creep, stress relaxation, buckling instabilities, and wave propagation [Stewart.2024, Alves.2021, Reese.1998, Golla.1985]. While these methods have been successful in many scenarios, they often require refined meshing strategies and explicit perturbations for specific problems. For example, in the instability analysis, traditional FEM always need artificial perturbations or imperfections which can lead to different buckling modes [Bazant.2010]. Moreover, in highly nonlinear and evolving systems, such as biological growth and morphogenesis, these conventional approaches can struggle to handle very complex geometries or require complex implementation strategies [BenAmar.2025, Liu.2024a]. The need for a more adaptive computational framework is evident.
@@
Figure 1: Viscoelastic materials and models. (a) Viscoelasticity is very common among different materials, e.g., hydrogels, bitumen [Edgeworth.1984], and soft biological tissues. (b-c) The viscoelastic models used for mechanical modeling. (b) The general Maxwell model. (c) The standard three-parameter model.
@@
Physics-informed neural network (PINN) has recently emerged as a promising alternative to solve some numerical problems [Raissi.2019, Review_2]. PINN offers a mesh-free approach that combines data and physical laws to solve differential equations by embedding the governing physics into the loss function of neural networks [Raissi.2019, Wang.2024]. It has been successfully applied to forward and inverse problems in fluid dynamics [Cai.2021, Zhu.2024], solid mechanics [Bai.2025, Zhuang.2021], and material science [Zhang.2022, Zhang.2024], demonstrating their capability to solve physical problems. For the viscoelastic research, the inverse viscoelastic problem has been studied using a deep neural network to obtain parametric and neural-network-based viscoelastic models from limited displacement data [Xu.2021]. A proposed ViscoelasticNet using PINN with appropriate constitutive models was applied to learn stress fields of viscoelastic fluids [Thakur.2024]. The deep energy method was developed to solve partial differential equations governing the deformation of hyperelastic and viscoelastic materials [Abueidda.2022]. It leverages the variational principles of mechanics, training networks to minimize a system’s potential energy functional. These pioneering applications validate the use of PINN and related deep learning techniques for mechanical simulation and viscoelastic constitutive modeling.
@@
For instability studies, buckling analysis of Kirchhoff plates has been implemented using the deep energy method with the buckling governing equations-related loss function to predict the critical buckling load [Zhuang.2021]. While this marks significant progress for elastic instabilities, the distinct phenomenon of creep buckling of viscoelastic structures—where instability develops over time under sustained load due to viscosity—necessitates a dedicated PINN formulation capable of handling long-term temporal integration and history-dependent material laws. Indeed, for viscoelastic structures, the creep buckling is one of the important failure modes, signifying failure after a period under steady load either via sudden snapping or excessive deflection [Miyazaki.2015]. Additionally, during growth and development of soft biological tissues with viscoelastic property, the symmetry broken and geometrical stability loss are important processes for tissue morphogenesis [Lin.2025, Moghe.2025, Harmansa.2023]. Despite the large potential of PINN, extending these approaches to capture creep buckling of viscoelastic structures, and to model the intricate interplay of mechanics and material evolution in viscoelastic growth and morphogenesis, presents difficulties and remains for investigation.
Addressing these challenges, this study develops and applies an energy-based PINN framework for modeling viscoelastic behavior. We extend the energy-based PINN framework [Bai.2025, Nguyen-Thanh.2020, Abueidda.2022, Zhuang.2021] to the setting of nonlinear viscoelasticity, where the governing equations depend not only on the instantaneous deformation but also on the full deformation history. This requires incorporating time-dependent evolution laws and coupling elastic and viscous energy contributions directly within the PINN formulation. Building on this foundation, we further examine how the dynamics of network optimization influence the numerical solution and show that optimizer-induced oscillations can serve as an intrinsic perturbation mechanism capable of initiating buckling in viscoelastic structures. In addition, we apply the framework to growth-induced deformation and morphogenesis, demonstrating that differential growth in cylindrical geometries can naturally lead to geometric instabilities and pattern formation. These results reveal that physics-informed neural networks can serve as an effective and flexible computational tool for simulating time-dependent material behavior and morphogenesis beyond the reach of conventional numerical methods. The remainder of this paper is organized as follows: Section 2 introduces the mathematical formulation of viscoelasticity. Section 3 details the implementation of PINN for viscoelastic simulation, and the benchmark of creep and stress relaxation. Section 4 explores the role of training oscillations in capturing creep buckling phenomena. Section 5 extends the framework to viscoelastic growth and morphogenesis, demonstrating its potential in biological applications. Finally, Section 6 discusses the advantages, limitations, and future research directions.
@@ 2.1 Governing Equation of Nonlinear Viscoelasticity
Let $\Omega$ be the configuration of a body in the three-dimensional Euclidean space $\mathbb{E}3$, and the corresponding initial configuration is $\Omega_{0}$. The deformation gradient tensor is $\mathbf{F}$, and $\mathbf{F}=\Delta\mathbf{u}+\mathbf{I}$, where $\mathbf{u}$ is the displacement vector and $\mathbf{I}$ is the identity tensor. For hyperelastic neo-Hookean materials, the stain energy density can be expressed as [Ogden.1997]
@@
where $\lambda$ is the Lamé’s first parameter, $G$ is shear modulus, $J=\text{det}\left(\mathbf{F}\right)$ is the Jacobian which measure the relative volume change, $I_{1}=B_{ii}$ is the first deformation invariant. $\mathbf{B}=\mathbf{F}\cdot\mathbf{F}^{T}$ is the left Cauchy-Green deformation tensor. The Cauchy stress formula can be deduced by differentiating the strain gradient density, as
@@
For large-deformation viscoelastic model, based on previous studies [Khajehsaeid.2014, Gamonpilas.2012], the constitutive equation can be written as
@@
where $m(t)$ is the dimensionless relaxation function and not changes during deformation. Using the general Maxwell model (Fig. 1b), $m(t)$ can be expressed by Prony series:
@@
where $\tau_{i}=\xi_{i}/E_{i}$ are relaxation times, $m_{\infty}+\sum_{i}m_{i}=1$. If we describe viscoelasticity by standard solid model (Fig. 1c), it has
@@ 2.2 Stress Relaxation and Creep
For viscoelastic stress relaxation condition, the deformation keeps constant, denoted as $\mathbf{F}_{0}$. Then, the Cauchy stress evolution equation can be obtained from Eq. (6), i.e.,
@@
where $\mathbf{B}_{0}=\mathbf{F}_{0}\cdot\mathbf{F}_{0}^{T}$, $J_{0}=\det\left(\mathbf{F}_{0}\right)$. Specially, under the small deformation and one-dimensional condition, the strain keeps constant as $\varepsilon_{0}$, and the stress evolution equation reduces to $\sigma(t)=E(t)\varepsilon_{0}$ [Christensen.1982]. When using the standard solid model Fig. 1c, the stress is
@@
The relaxation time $\tau$ means the needed time that the material relaxed all relatable stresses if the stress is relaxed with the initial rate $\dot{\sigma}(0)=-(E_{1}\varepsilon_{0})/{\tau}$.
Creep is defined as the tendency of a solid material to deform permanently under the influence of constant stress. For creep condition, the stress keeps constant, i.e., $\sigma_{0}$. The strain creep function can be written as $\mathbf{e}(t)=J(t)\sigma_{0}$, where $J(t)$ is the creep compliance function and $\mathbf{e}$ is the strain tensor. The creep compliance can be obtained by applying the Laplace transform and inverting the relaxation function. It may be not solved analytically under large deformation, but the creep behavior can be captured easily by numerical method, like finite element analysis. Specially, for the small deformed one-dimensional case, the creep process becomes using the standard solid model (Fig. 1c)
@@
where the retardation time $\tau_{r}$ is defined as $\tau_{r}=\tau(E_{\infty}+E_{1})/E_{\infty}$, which can characterized the creep behavior. The creep rate is
@@
If the material creep with the initial creep rate $\dot{\varepsilon}(0)={\sigma_{0}E_{1}}/[{\tau_{r}E_{\infty}(E_{\infty}+E_{1})}]$, the duration of the creep process from $\varepsilon(0)=\sigma_{0}/(E_{1}+E_{\infty})$ to $\varepsilon(\infty)=\sigma_{0}/E_{\infty}$ is $\tau_{r}$.
@@ 2.3 Creep Buckling of Viscoelastic Structures
Creep buckling refers to a specific type of structural failure that occurs in materials subjected to a sustained compressive load over an extended period [Miyazaki.2015, Liu.2022]. As the material creeps, its effective rigidity decreases. This reduction can lower the critical buckling load, i.e., the maximum load that a structure can withstand before it buckles, leading to premature failure even when loads are lower than the instantaneous buckling thresholds.
@@
where $l$ is the length of cantilever, $E$ is the Young’s modulus of the material and $I_{g}$ is the area moment of inertia. For the cross-section with area $A$, the critical pressure is
@@
For a linear elastic thin-walled cylindrical shell under uniform radial pressure, the critical buckling pressure can be estimated by [Miyazaki.2015]
@@
where $d$ is the cylinder diameter, and $a$ is the wall thickness. If a structure, such as a slender beam or a thin-walled cylindrical shell, has viscoelastic properties, its stiffness will become smaller and the deformation will become larger under the constant external load with time. Thus, its critical buckling force will be smaller than its original state. For creep buckling analysis, we can easily use the quasi-static approach, that is, buckling occurs at time $t_{cr}$ when the applied pressure $p$ equals the critical elastic buckling load calculated using the current relaxed modulus $E(t_{cr})$ and Eqs. (13) or (14).
@@ 2.4 Viscoelastic Tissue Growth
Viscoelasticity is a fundamental property of biological soft tissues, characterized by their ability to exhibit both elastic and viscous behavior when subjected to deformation [Chaudhuri.2020, Clement.2017]. The stress modulated growth models have been developed [Walker.2023, Sun.2022, Huang.2024] since the stress-growth law proposed by Fung [Fung.1990]. In this kind of growth model, the deformation gradient tensor is decomposed as (Fig. 2)
@@
where $\mathbf{F}_{a}$ is the mechanical deformation tensor and $\mathbf{F}_{g}$ is the growth tensor. The growth tensor is always written as $\mathbf{F}_{g}=g_{i}\mathbf{I}$, where $g_{i}$ is the growth ratio. The growth rate is determined by the stress state [Huang.2024, Lin.2025], i.e.,
@@
where $\mathbf{b}^{g}=b^{g}_{i}\mathbf{I}$ is the biochemical forces driving growth [Buskohl.2014], and $k$ is a constant to regulate the stress effect on growth. For viscoelastic tissues, the stress $\sigma$ is obtained by Eq. (6) using deformation gradient tensor $\mathbf{F}_{a}$, which can relax as time goes. Stress relaxation slows residual stress accumulation, leading to increased growth rates [Lin.2025].
@@
A critical consequence of the differential or constrained growth described by the growth tensor $\mathbf{F}_{g}$, coupled with the viscoelastic response, is the generation of internal residual stresses [Erlich.2024, Lin.2025]. When growth is non-uniform (e.g., faster surface growth relative to the bulk) or spatially constrained, significant compressive stresses can accumulate within the tissue over time. If these growth-induced compressive stresses reach a critical threshold, the initial configuration of the tissue may lose stability, leading to a buckling event [Li.2011, Huang.2024]. This mechanical instability provides a fundamental mechanism for morphogenesis, driving the spontaneous formation of complex patterns such as wrinkles, folds, or creases observed during development and other biological processes.